OpenStudy (anonymous):

Maya and Sally were the two finalists in a singing competition. The person with the most votes from the audience was chosen as the winner. Sally received 25% of the total votes and lost by 62,500 votes. No person in the audience was allowed to vote more than once. The equation below represents this situation, where y is the total number of votes cast and x is the number of votes Maya received. : x - 62,500 = 0.25y If a total of 125,000 people voted, how many votes did Maya receive? Do not enter comma for placeholder values. Numerical Answers Expected!

OpenStudy (anonymous):

sally recieved 25% of total votes total votes =125000 so y=125000*25/100 y=31250 so total -y=x =93750

OpenStudy (anonymous):

As the number of pages in a photo book increases, the price of the book also increases. There is an additional shipping charge of 15%. The price of a book can be modeled by the equation below, where P = the price of the book, 20 is the printing charge, 0.5 is the charge per page, and x = the number of pages. : P = (20 + 0.5x) + 0.15(20 + 0.5x) Jennifer wants to purchase a book but only has $62.10 to spend. What is the maximum number of pages she can have in her book? x = ______________________ pages

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

yes hold on

OpenStudy (anonymous):

his time you have to plug in the 62.10 to the P and solve for x. Let me see if I can help. I know that the new equation is now 62.10=(20+0.5x)+0.15(20+0.5x) so, x would equal 68, the number of pages then

OpenStudy (anonymous):

one more pretty plz

OpenStudy (anonymous):

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below, where A = the pounds of Arabica and R = pounds of Robusta. : 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

then it says A=1000 so now the equation is 700×1000+1200R=1,000,000 or 700,000+1200R=1,000,000 which i would solve in two steps 1) subtract 7000,000 to get 1200R=300,000 2) then divide by 1200 to get R=250